Find the inverse of the given function.

h(x)=sin-1 sqrt x


1
Expert's answer
2021-10-13T15:10:22-0400

h(x)= "sin^{-1}(\sqrt{x})"

Inverse of h(x) will exist if it is bijective function.

So Domain of h(x) is [0,1] and range is [0,"\frac{π}{2}]"

Let y = "sin^{-1}(\sqrt{x})"

=> Sin(y) = "\sqrt{x}"

=> x = sin²y

So inverse of h(x) , h"^{-1}(x)=" sin²(x) with domain [0,"\frac{π}{2}]" and range [0,1]